# FakeScope: Large Multimodal Expert Model for Transparent AI-Generated Image Forensics

### v1

@@ Abstract
The rapid and unrestrained advancement of generative artificial intelligence (AI) presents a double-edged sword: while enabling unprecedented creativity, it also facilitates the generation of highly convincing deceptive content, undermining societal trust. As image generation techniques become increasingly sophisticated, detecting synthetic images is no longer just a binary task—it necessitates interpretable, context-aware methodologies that enhance trustworthiness and transparency. However, existing detection models primarily focus on classification, offering limited explanatory insights into image authenticity. In this work, we propose FakeScope, an expert multimodal model (LMM) tailored for AI-generated image forensics, which not only identifies AI-synthetic images with high accuracy but also provides rich, interpretable, and query-driven forensic insights. To this end, we first construct FakeChain dataset that contains linguistic authenticity reasoning based on visual trace evidence, developed through a novel human-machine collaborative framework. Building upon it, we further present FakeInstruct, the largest multimodal instruction tuning dataset containing 2 million visual instructions tailored to enhance forensic awareness in LMMs. Leveraging the knowledge of FakeInstruct, FakeScope achieves state-of-the-art performance in both closed-ended and open-ended forensic scenarios. It can distinguish synthetic images with high accuracy while offering coherent and insightful explanations, free-form discussions on fine-grained forgery attributes, and actionable enhancement strategies. Notably, despite being trained exclusively on qualitative hard labels, FakeScope demonstrates remarkable zero-shot quantitative capability on detection, enabled by our proposed token-based probability estimation strategy. Furthermore, FakeScope exhibits strong generalization and in-the-wild ability, ensuring its applicability in real-world scenarios. The data, model, and demo will be publicly released on https://github.com/Yixuan423/FakeScope.
Index Terms:
AI-generated image detection, large multimodal models, forensic investigation, trustworthiness.
@@ I Introduction
With the rapid development of image generation techniques, AI-synthesized images are becoming increasingly indistinguishable from genuine ones [1, 2]. While this prosperity has driven innovation, it also introduces significant cybersecurity threats, including misinformation, propaganda, and fraud. In response, substantial efforts have been made to detect AI-generated content in recent years [3, 4, 5, 6]. However, as generative models continue to evolve, concerns over the reliability and trustworthiness of detection systems in real-world applications persist. Addressing these challenges requires enhancing the transparency of AI-generated image detection models. We advocate for a paradigm shift in which image synthesis detection goes beyond binary classification, incorporating explainability to provide contextual insights into authenticity assisting trustworthiness.
@@
Figure 1: Comparative illustration of image forensic model evolution. (a) Binary detectors; (b) general LMMs; and (c) FakeScope, which translates visual abnormalities into human-understandable cues—much like trace evidence in forensics—using natural language.
@@
Figure 2: Contributions of this work. (a) FakeChain dataset, containing reasoning on image authenticity, constructed via the proposed ACoTI strategy (Sec. III); (b) FakeInstruct dataset, containing 2 million visual instructions of image forensic knowledge (Sec. IV); (c) FakeScope model, the expert model for transparent AI-generated image forensics, capable of multi-dimensional forensic capabilities (Sec. V).
@@
Until recently, AI-generated image detection was predominantly framed as a binary classification task. Conventional approaches rely on unimodal architectures such as convolutional neural networks (CNNs) and vision transformers (ViTs) to classify images as either real or synthetic, significantly aiding in mitigating the spread of synthetic content. However, these early methods largely overlooked the need for transparency, offering little more than abstract heatmaps derived from deep neural network features as visual explanations. Recent research has shown that AI-generated images increasingly pass sanity checks by humans and machines [7, 1], leading to human-machine or machine-machine disagreements. As illustrated in Fig. 1 (a), different detectors can yield inconsistent decisions, significantly reducing practical utility. In contrast, in Fig.1(c), a transparent model can provide additional insights via visual trace evidence to enhance interpretability. Therefore, we anticipate more trustworthy paradigms that exceed binary detectors for achieving transparent forensics of AI-generated images. Such a model would enable human oversight, mitigate detection biases, and foster better confidence in AI-generated content forensics [2].
@@
With the advent of multimodal models, visual-language approaches have started to emerge. Zhang et al. [8] introduced textual template descriptions to improve explainable deepfake face detection. Sun et al. [9] formulated brief descriptions by integrating forgery regions and candidate types using a predefined template. Li et al. [7] and Ye et al. [10] explored the capabilities of large multimodal models (LMMs) in the forensic analysis of AI-generated images. In [11, 12, 13], LMMs are incorporated to localize tampering forgeries and exhibits substantial effectiveness. Besides, LMMs are employed to assist in improving detection accuracy. In [14] and [15], conventional detectors are combined with LMMs to improve detection accuracy. Wen et al. [16] train a binary classifier with ViT intermediate features and use LMMs to explain the detection. These studies establish a concrete foundation for the transition from unimodal to multimodal approaches in the area of AI-generated image forensics. Nevertheless, due to the inherent limitations of LMMs as classifiers [17], previous approaches have struggled to balance detection accuracy and transparency, often relying on separate strategies and training procedures to fulfill multi-dimensional forensic tasks. Therefore, there is an urgent need for a unified forensic expert model for AI-generated images capable of multiple forensic tasks in diversified scenarios.
@@
Inspired by the above insights, we present FakeScope, a unified expert model for transparent and versatile AI-generated image forensics. It leverages the remarkable cross-modality ability of generally pretrained LMMs while being systematically enhanced for universal forensic awareness. Specifically, it is designed with three key objectives: (1) comprehension on trace evidence and authenticity; (2) faithful response to diverse user queries; and (3) qualitative and quantitative detection without model extension. To achieve these, we collect and leverage forensic knowledge to improve LMMs’ forensic awareness on AI-generated images. As shown in Fig. 2(a), we first construct FakeChain, a large-scale multimodal dataset containing long and comprehensive reasoning on image authenticity as forensic knowledge source, developed via a cost-efficient human-in-the-loop strategy. We further enrich FakeChain to create FakeInstruct, a gigantic-scale dataset comprising 2 million visual instructions to enhance diverse forensic abilities of LMMs (Fig. 2(b)). Extensive experiments demonstrate that FakeInstruct consistently improves the forensic capabilities of general LMMs, and the proposed FakeScope (Fig. 2(c)) exhibits state-of-the-art (SOTA) performance in detecting, explaining, analyzing, and discussing AI-generated images. In particular, it can generalize well to unseen image content, generation models, and even user-created contents. Our core contributions are as follows.
@@
  • We present FakeChain, a large-scale multimodal dataset with coherent reasoning flow from visual trace evidence to image authenticity. It is collected through our proposed Anthropomorphic Chain-of-Thought Inference (ACoTI) scheme, a cost-effective human-machine collaboration strategy that requires minimal human annotation while ensuring reliability.
@@
  • We develop FakeInstruct, the first million-scale dataset designed to enhance LMMs’ multi-dimensional forensic capabilities in AI-generated image detection.
@@
  • We propose FakeScope, a multimodal expert model for transparent AI-generated image detection that integrates both detection and query-based forensic analysis. We further propose a zero-shot token-based strategy for estimating detection probabilities. Extensive experiments validate FakeScope’s superiority in both detection and transparency-related tasks, outperforming other LMMs and binary models across diverse datasets.
II Related Works
Recently, the rapid advancement of image generation technology has driven growing interest in AI-generated image forensics, and the evolution of LMMs has opened new avenues for enhancing the trustworthiness of image forensic systems. We first outline the progression of AI-generated image detection techniques, followed by an overview of the current developments in LMMs.
@@ II-A AI-generated Image Detection
The detection of AI-generated images aims to differentiate fake images (synthesized by generative AI algorithms) from real ones (created or captured by humans). This problem is commonly formulated as a binary classification task. Preexistent approaches generally incorporate forensic patterns to classify, which can be categorized into the flaw-based [18, 19] and fingerprints-based [20, 21, 22, 23] according to their forensic ground. The flaw-based methods capture the intricate details in low (e.g., texture [18]) or high (e.g., physical rules [19, 24]) semantic levels, while the fingerprints-based detectors identify the invisible leftover patterns of generators (e.g., frequency artifacts [20, 21, 22, 23]). Moreover, investigations on model’s generalization ability towards unseen  [25, 26, 4, 3, 5, 6] and generator attribution [27, 28] have emerged. Some of them assume to be only accessible to real images and utilize generators (like an autoencoder) to simulate generative artifacts [29, 22, 26, 25]. Although remarkable, current fake image detection models still fall short in the interpretability towards human [30], which causes the detection results hard-to-explain for the general populace and weakens the credibility of model output. Therefore, developing an expert dedicated to fake image detection with both high detection accuracy and transparency is currently urgently needed.
@@ II-B Large Multimodal Models
Typical LMMs compromise a modality encoder, a large language model (LLM), with a modality interface for cross-modal connection [31, 32]. The specialty of LMM lies in their ability to effectively receive, process, and generate multimodal information. Currently, LMMs are developing by leaps and bounds and have exhibited remarkable visual-language capabilities on par with or even surpassing the human, including the closed-source models (e.g., GPT-4V [33], Gemini [34], and Claude [35]) and the open-source ones (e.g., LLaVA [36], mPLUG-Owl [37], deepseek [38]). Furthermore, to adapt general-purpose LMMs to specialized fields, recent researchers proposed various expert models based on visual instruction tuning  [39, 40, 41, 36]. Despite great progress, recent studies [7] indicate that LMMs’ forensic capabilities are still unsatisfactory, especially when it comes to identifying fine-grained forgery aspects. On the other hand, human crowd-sourcing data for visual instruction tuning becomes increasingly hard to get, partially due to the process being time-consuming, especially when multimodal information is considered. In this work, to infuse the forensic abilities on AI-generated images into foundational LMMs, we construct the first visual instruction tuning dataset centering on fake image detection, the FakeInstruct, aiming at a unified multimodal paradigm based on LMMs for more transparent image forensics.
@@ III FakeChain: Reasoning on Image Authenticity
In this section, we elaborate on the construction of FakeChain, a large-scale dataset with long and comprehensive reasoning on underlying trace evidence for authenticity. It functions as the source of forensic knowledge and further contributes to building the instruction dataset and the forensic model. For construction, we first prepare images (Sec. III-A) and then collect authenticity reasoning. While human annotation remains the most intuitive and widely applied approach for training LMMs [42, 36], acquiring sufficiently fine-grained dataset for authenticity reasoning is highly labor-intensive. To mitigate the prohibitive costs, a natural idea is to use a human-machine collaboration strategy. Therefore, we first collect human expert feedback (Sec. III-B) as weak supervision in the proposed Anthropomorphic Chain-of-Thought Inference (ACoTI) strategy (Sec. III-C) for forensic reasoning collection.
@@ III-A Image Preparation
FakeChain contains both AI-generated and real images in equal quantities. In contrast to the one-to-all generalization setting in some works [25, 26, 4, 3, 5, 6], FakeChain is deliberately constructed with a high degree of diversity to capture a broad range of generative patterns. In particular, AI-generated images function as the knowledge source of forensic trace evidence. To ensure content diversity, we sampled 23,797 AI-generated images from multiple sources. These generators span 17 types, including GAN-based, diffusion-based, auto-regressive, and proprietary models. In specific, we choose representative ProGAN [43], StyleGAN [44], BigGAN [45], CogView2 [46], ADM [47], IF [48], FuseDream [49], VQDM [50], Glide [51], SD [52], SDXL [53], FLUX.1 [54], Firefly [55], Dalle2 [56], Dalle3 [57], Wukong [58] and Midjourney [59]. Also, we collect 23,797 real images from DIV2K [60], RAISE [61], and ImageNet [62]. The details of image components are provided in the supplementary material (SM).
@@ III-B Human Expert Feedback
As shown in Fig. 3, we require a few human annotations in step1-steer and step2-demonstrate of ACoTI, including the trace evidence category and reasoning demonstration respectively. Therefore, we conduct a well-controlled in-lab experiment involving 30 trained human experts with experience in image generation as annotators, ensuring qualified understanding of generative patterns. For step1-steer, experts are required to correctly distinguish whether the image is AI-generated, and then select supporting categories, ensuring only accurate judgments contribute to the dataset. In practice, we provide a category set for reference $\mathcal{F}$ = {texture, edge, clarity, light&shadow, anatomy, layout, symmetry, reflection, perspective, physics, shape, theme, content deficiency, distortion, unrealistic, overall hue[7], which can characterize authenticity across different levels of trace evidence. Each image is annotated by two and only the intersection of selected categories is retained to improve reliability. Additionally, experts can provide supplementary authenticity-related information at will, such as specific area of occurrence and commonsense knowledge. For step 2-demonstrate, we fix 50 fake and 50 genuine images, and employ experts to annotate manually with detailed reasoning. These expert-annotated images constitute the exemplar sets $\{{\chi}^{+},{\chi}^{-}\}$ as illustrated in upper Fig. 3, where + and - denote AI-generated and genuine images respectively. To ensure balance, we manually guarantee each category of $\mathcal{F}$ emerges in at least three exemplars. Herein, we establish the necessary human supervision for the subsequent reasoning collection.

### v2

@@ Abstract
The rapid and unrestrained advancement of generative artificial intelligence (AI) presents a double-edged sword. While enabling unprecedented creativity, it also facilitates the generation of highly convincing content, undermining societal trust. As image generation techniques become increasingly sophisticated, detecting synthetic images is no longer just a binary task—it necessitates explainable methodologies to enhance trustworthiness and transparency. However, existing detection models primarily focus on classification, offering limited explanatory insights. To address these limitations, we propose FakeScope, an expert large multimodal model (LMM) tailored for AI-generated image forensics, which not only identifies synthetic images with high accuracy but also delivers rich query-contingent forensic insights. At the foundation of our approach is FakeChain, a large-scale dataset containing structured forensic reasoning based on visual trace evidence, constructed via a human-machine collaborative framework. Then we develop FakeInstruct, the largest multimodal instruction tuning dataset to date, comprising two million visual instructions that instill nuanced forensic awareness into LMMs. Empowered by FakeInstruct, FakeScope achieves state-of-the-art performance in both closed-ended and open-ended forensic scenarios. It can accurately distinguish synthetic images, provide coherent explanations, discuss fine-grained forgery artifacts, and suggest actionable enhancement strategies. Notably, despite being trained exclusively on qualitative hard labels, FakeScope demonstrates remarkable zero-shot quantitative detection capability via our proposed token-based probability estimation strategy. Furthermore, it shows robust generalization across unseen image generators and performs reliably under in-the-wild scenarios. The data, model, and demo will be publicly released on https://github.com/Yixuan423/FakeScope.
@@ I Introduction
With the rapid development of image generation techniques, AI-synthesized images are becoming increasingly indistinguishable from genuine ones [76, 85]. While this prosperity has driven innovation, it also introduces significant cybersecurity risks, including misinformation, propaganda, and fraud. In response, substantial efforts have been made to detect AI-generated content in recent years [53, 12, 13, 60]. However, as generative models continue to evolve, concerns over the reliability and trustworthiness of detection systems in real-world applications persist. Addressing these challenges requires enhancing the transparency of AI-generated image detection models. We advocate for the potential in which image synthesis detection goes beyond binary classification, incorporating explainability to provide contextual insights into authenticity, assisting trustworthiness.
@@
Figure 1: Comparative illustration of image forensic model evolution. (a) Binary detectors; (b) general LMMs; and (c) FakeScope, which translates visual abnormalities into human-understandable cues, much like trace evidence in forensics, using natural language.
@@
Figure 2: Contributions of this work. (a) FakeChain dataset, containing reasoning on image authenticity, constructed via the proposed ACoTI strategy (Sec. III); (b) FakeInstruct dataset, containing two million visual instructions concerning image forensic knowledge (Sec. IV); (c) FakeScope model, the expert model for transparent AI-generated image forensics, capable of multi-dimensional forensic capabilities (Sec. V).
@@
Until recently, AI-generated image detection was predominantly framed as a binary classification task. Conventional approaches rely on unimodal architectures such as convolutional neural networks (CNNs) and vision transformers (ViTs) to classify images as either real or synthetic, significantly aiding in mitigating the spread of synthetic content [68, 22, 24]. However, these early methods largely overlooked the need for transparency, offering little more than abstract heatmaps derived from deep neural network features as visual explanations. Recent research has shown that AI-generated images increasingly pass sanity checks by humans and machines [40, 76], leading to human-machine or machine-machine disagreements. As illustrated in Fig. 1(a), different detectors can yield inconsistent decisions, significantly reducing practical utility. In contrast, in Fig.1(c), a transparent [^1] Following [28, 44], transparency herein refers to a composite spectrum including explainability, traceability, and justifiability. forensic model can go beyond binary decisions by offering visual-evidence-based justifications, supporting query-driven forensic analysis, and estimating authenticity probabilities in a unified and interpretable manner. Such capabilities are key to enabling human oversight, mitigate detection biases, and foster better confidence in AI-generated content forensics [28, 85].
@@
With the advent of multimodal models, vision-language approaches have started to emerge [66, 31, 80, 9, 29, 71, 84]. Zhang et al. [80] introduced textual template descriptions to improve explainable deepfake face detection. Sun et al. [66] formulated brief descriptions by integrating forgery regions and candidate types using a predefined template. Li et al. [40] and Ye et al. [77] explored the capabilities of large multimodal models (LMMs) in the forensic analysis of AI-generated images. In [75, 46, 32], LMMs are incorporated to localize tampering forgeries and exhibit substantial effectiveness. Besides, LMMs are employed to assist in improving detection accuracy. In [31] and [9], conventional detectors are combined with LMMs to improve detection accuracy. Wen et al. [71] train a binary classifier with ViT intermediate features and use LMMs to explain the detection result. These studies have established a concrete foundation for transitioning from unimodal to multimodal approaches in AI-generated image forensics. Nevertheless, due to the inherent limitations of LMMs as classifiers [81], previous approaches have struggled to achieve high detection performance and transparency simultaneously. Most existing works rely on separate strategies and training procedures, such as introducing additional classifier heads or independently trained explanation modules, to fulfill different forensic demands. Such fragmented pipeline and weak coupling can lead to disconnected probabilistic reasoning space across tasks and hindering the model from reusing shared forensic knowledge, further constraining the model’s potential. This motivates the need for a unified forensic expert model for AI-generated images capable of performing multiple forensic tasks in diversified scenarios.
@@
Inspired by the insights, we present FakeScope, a unified expert model for transparent and versatile AI-generated image forensics. It leverages the remarkable cross-modality ability of generally pretrained LMMs while being systematically enhanced for forensic awareness. Specifically, it is designed with three key objectives: (1) comprehension of visual trace evidence and authenticity; (2) faithful response to diverse user queries; and (3) qualitative and quantitative detection without model extension. To achieve these, we collect and leverage domain-specific knowledge to improve LMMs’ forensic awareness of AI-generated images. Our emphasis is therefore on output-level image forensic transparency rather than on modeling the causal interpretability of internal LMM mechanisms. As shown in Fig. 2(a), we first construct FakeChain, a large-scale multimodal dataset containing long and comprehensive reasoning on image authenticity as a forensic knowledge source, developed via a cost-efficient human-in-the-loop strategy. We further enrich FakeChain to create FakeInstruct, a dataset comprising two million visual instructions to enhance the diverse forensic abilities of LMMs (Fig. 2(b)). Extensive experiments demonstrate that FakeInstruct consistently improves the forensic capabilities of general LMMs, and the proposed FakeScope (Fig. 2(c)) exhibits state-of-the-art (SOTA) performance in detecting, explaining, analyzing, and discussing AI-generated images. In particular, it generalizes well to unseen image content, generation models, and user-created content. Our core contributions are as follows.
@@
  • We present FakeChain, a large-scale multimodal dataset with structured reasoning from visual trace evidence to image authenticity. It is built with our proposed Anthropomorphic Chain-of-Thought Inference (ACoTI) scheme. This cost-effective human-machine collaboration strategy requires minimal human annotation while ensuring data reliability.
@@
  • We develop FakeInstruct, the first million-scale visual instruction dataset for enhancing LMMs’ forensic capabilities in AI-generated image detection, exhibiting consistent enhancement to diverse base models.
@@
  • We propose FakeScope, a multimodal forensic expert built on FakeInstruct for transparent AI-generated image detection. It unifies detection and query-based forensic analysis, and further incorporates a zero-shot token-based strategy for probability estimation. Extensive experiments validate its superiority in detection and transparency-related tasks, outperforming other LMMs and binary models across diverse datasets.
II Related Work
In this section, we first outline the progression of AI-generated image detection techniques, followed by an overview of recent developments in LMMs.
@@ II-A AI-generated Image Detection
The detection of AI-generated images aims to differentiate fake images (synthesized by generative AI algorithms) from real ones (created or captured by humans). This problem is commonly formulated as a binary classification task. Preexisting approaches generally incorporate forensic patterns to classify, which can be categorized into the flaw-based [83, 11] and fingerprints-based [68, 22, 24, 72] according to their forensic ground. The flaw-based methods capture the intricate details in low (e.g., texture [83]) or high (e.g., physical rules [11, 63]) semantic levels. At the same time, the fingerprints-based detectors identify the invisible leftover patterns of generators (e.g., frequency artifacts [68, 22, 24, 72]). Moreover, investigations on model’s generalization ability towards unseen  [43, 33, 12, 53, 13, 60] and generator attribution [64, 69] have emerged. Some of them are assumed to be only accessible to real images and utilize generators (like an autoencoder) to simulate generative artifacts [79, 24, 33, 43]. Although remarkable, current fake image detection models still fall short in interpretability towards humans [41], which makes the detection results hard to explain for the general populace and weakens the credibility of the model output. Therefore, developing an expert dedicated to fake image detection with high detection accuracy and transparency is urgently needed.
@@ II-B Large Multimodal Models
Typical LMMs comprise a modality encoder, a large language model (LLM), with a modality interface for cross-modal connection [20, 30]. The specialty of LMMs lies in their ability to effectively receive, process, and generate multimodal information. Currently, LMMs are developing by leaps and bounds and have exhibited remarkable visual-language capabilities on par with or even surpassing the human, including the closed-source models (e.g., GPT-4 [54], Gemini [67], and Claude [3]) and the open-source ones (e.g., LLaVA [45], mPLUG-Owl [78], Deepseek [49]). Furthermore, to adapt general-purpose LMMs to specialized fields, recent researchers proposed various expert models based on visual instruction tuning  [48, 50, 21, 45]. Despite great progress, recent studies [40] indicate that LMMs’ forensic capabilities are still unsatisfactory, especially when identifying fine-grained forgery aspects. In this work, to infuse the forensic abilities of AI-generated images into foundational LMMs, we construct the first visual instruction tuning dataset centering on fake image detection, the FakeInstruct, aiming at a unified multimodal paradigm based on LMMs for a more transparent image forensics.
@@ III FakeChain: Reasoning on Image Authenticity
In this section, we elaborate on the construction of FakeChain, a large-scale dataset of visual trace-based authenticity reasoning. FakeChain serves as the primary source of forensic knowledge and further supports the construction of the instruction dataset and the expert model.
We first prepare images (Sec. III-A) and then collect authenticity reasoning. While human annotation remains the most intuitive approach for training LMMs [73, 45], collecting fine-grained authenticity reasoning at scale is highly labor-intensive. To reduce this cost, we adopt a human-machine collaborative strategy (Fig. 3) that uses limited human annotations as weak supervision in three steps:
1. Steer, which guides reasoning with human-annotated trace-evidence categories;
2. Demonstrate, which provides human reasoning exemplars;
3. Enlighten which elicits visual reasoning from strong models through few-shot chain-of-thought prompting.
Building on this scheme, we first collect human expert feedback (Sec. III-B) and propose Anthropomorphic Chain-of-Thought Inference (ACoTI) strategy (Sec. III-C2) to enlighten strong models under the steer of a few human annotations assisted by human reasoning demonstrations. With ACoTI, we construct the FakeChain dataset. Details are elaborated below.
@@ III-A Image Preparation
FakeChain contains a equal number of AI-generated images and real images, totaling 47,594. Unlike the one-to-all generalization setting in [43, 33, 12, 53, 13, 60], FakeChain is deliberately designed for high diversity to cover a wide range of generative patterns. In particular, AI-generated images function as the knowledge source of forensic trace evidence and are sampled from multiple sources. The generators span 17 types, including GAN-based, diffusion-based, auto-regressive, and proprietary models. Specifically, we choose representative ProGAN [35], StyleGAN [36], BigGAN [6], CogView2 [19], ADM [18], IF [16], FuseDream [47], VQDM [27], Glide [52], SD [61], SDXL [56], FLUX.1 [23], Firefly [1], Dalle2 [58], Dalle3 [65], Wukong [74] and Midjourney [51]. Also, we collect 23,797 real images from DIV2K [2], RAISE [15], and ImageNet [17], ensuring content diversity. The details are provided in the supplementary material (SM).
@@ III-B Human Expert Feedback
As shown in Fig. 3, we require a few human annotations in step1-steer and step2-demonstrate, including the trace evidence category and reasoning demonstration. Therefore, we conduct a controlled in-lab experiment involving 30 trained human experts experienced in image generation, ensuring qualified understanding of generative patterns. In step1-steer, experts first distinguish whether an image is AI-generated and then select supporting categories. This ensures only correct annotations contribute to the dataset. In practice, we provide a category set for reference $\mathcal{F}$ = {texture, edge, clarity, light&shadow, anatomy, layout, symmetry, reflection, perspective, physics, shape, theme, content deficiency, distortion, unrealistic, overall hue[40], covering multiple levels of forensic evidence. To avoid overly broad selections, annotators are instructed to choose only verifiable cues rather than all plausible categories. Practically, each image receives at most six categories. Every image is annotated by two experts, and only the intersection of their selections is kept to reduce annotator-specific bias. Additionally, experts can provide authenticity-related observations at will, such as specific affected regions or relevant commonsense cues. In step 2-demonstrate, we fix 50 fake and 50 genuine images and employ experts to annotate manually with detailed reasoning. These annotations constitute the exemplar set $\{{\chi}^{+},{\chi}^{-}\}$ shown in upper Fig. 3, where + and - denote AI-generated and genuine images respectively. To ensure coverage, we manually guarantee that each category in $\mathcal{F}$ emerges in at least three exemplars. This process establishes the human supervision needed for subsequent reasoning collection.
Figure 4: The composition of the FakeInstruct, which is derived from the FakeChain dataset, containing 47K visual instructions on absolute authenticity judgments, 95K instructions on holistic reasoning, 715K on fine-grained visual trace evidence, and 1190K on extended knowledge and discussions. The million-scale diversified visual instructions of FakeInstruct enable LMMs with a broad and fine-grained understanding of image authenticity, ensuring LMMs are equipped to handle diverse forensic tasks with both interpretability and accuracy.